# The impact of future D and B meson measurements with the SMOG2 program at LHCb on the determination of nuclear parton distribution functions

### v1

@@ Abstract
We perform an analysis of the potential impact of future fixed-target data, to be measured by the SMOG2 experiment by using various nuclear targets in the LHCb detector, on nuclear parton distribution functions. The analysis is performed by using the reweighting method, that allow us to investigate the impact of future $D$- and $B$-meson measurements at SMOG2 on different extractions of nuclear parton densities. We show how these future data can reduce the uncertainty on the gluon and the sea quarks distributions in nuclei.
I Introduction
@@
One of the crucial aspects of high-energy nuclear physics is studying the structure of matter, in particular, in terms of its smallest components, quarks and gluons, collectively referred to as partons. Within the collinear factorisation approach Bodwin:1984hc; Collins:1985ue, the partonic structure of a hadron is described by parton distribution functions (PDFs), which represent the probability of finding partons with a given fraction $x$ of the proton longitudinal momentum. PDFs encapsulate the non-perturbative information about hadron content, and are determined through fits to experimental data Ethier:2020way; Gao:2017yyd; Kovarik:2019xvh; Klasen:2023uqj.
@@
Furthermore, the partonic structure of nucleons bound in a nucleus differs from that of a free proton. This phenomenon was first identified as the EMC effect EuropeanMuon:1981gyc. Subsequent studies have shown that this modification varies with $x$ and we can now recognise four characteristic regions of nuclear modifications EuropeanMuon:1988lbf; EuropeanMuon:1988tpw; EuropeanMuon:1983wih; Goodman:1981hc; Bodek:1983ec; BCDMS:1985dor; Klein:2019qfb. These are identified as: (i) the shadowing region at $x\lesssim 0.1$ (characterised by a depletion of nuclear cross-sections compared to the free-nucleon ones), (ii) the antishadowing region at $0.25\lesssim x\lesssim 0.1$ (characterised by a corresponding enhancement), (iii) the EMC region $0.75\lesssim x\lesssim 0.25$ (characterised by depletion), and (iv) the Fermi-motion region $x\gtrsim 0.75$ (characterised by enhancement). The strength of these effects also changes with the atomic mass number of the nucleus.
@@
Factorisation framework directly links these modifications observed on the level of cross sections with the behaviour of nuclear PDFs (nPDFs). Therefore, a precise determination of nPDFs is essential for using perturbative quantum chromodynamics (pQCD) to interpret experimental data from high-energy nuclear reactions. For example, this information is crucial for using charm and bottom mesons to study the hot and dense partonic matter created in ultrarelativistic heavy-ion collisions at the Relativistic Heavy Ion Collider at Brookhaven National Laboratory STAR:2018zdy or at the Large Hadron Collider (LHC) at CERN (for example ALICE:2022wpn), as well as for computing cross sections in astrophysics, e.g. Bhattacharya:2016jce. Despite significant progress over the past decades, our understanding of nPDFs, especially the gluon ones, remains incomplete. For instance, there is a limited knowledge about how nuclear effects vary with the number of nucleons in a nucleus.
@@
Charm and bottom quarks in high-energy reactions are predominantly produced via gluon fusion. Therefore, the yields of charm and bottom hadrons can be used to determine gluon nPDFs Kusina:2017gkz; Kusina:2020dki. The LHCb experiment facilitates an effective measurement of $D$ and $B$ mesons in proton-nucleus collisions across various ion species making it a perfect place for robust study on nuclear gluon PDFs.
@@
In this paper, we review the prospects for investigating nuclear parton distribution functions using $p$A pseudo-data at $\sqrt{s_{NN}}=113$ GeV from the SMOG2 fixed-target program at the LHCb experiment BoenteGarcia:2024kba. The pseudo-data used for the analysis are generated with MadGraph5_aMC@NLO Flore:2025ync. The paper is organised as follows: in Section II we begin with a brief overview of the main features of fixed-target collisions using LHC beams. Next, we describe the experimental setup for these collisions at the LHCb experiment, including the ion species that can be used and the expected parameters of this experimental program. Section III introduces the formalism employed to evaluate the impact of the anticipated LHCb data on nPDFs. Finally, we present our results in Section IV and summarise our findings in Section V.
II Fixed target program at the LHCb experiment
II.1 Overview of the fixed-target program at the LHC
@@
An experiment utilising the LHC beams and a fixed target offers numerous, exciting and timely opportunities Hadjidakis:2018ifr; Brodsky:2012vg. The high energy and intensity of the proton and ion beams at the LHC enable compelling studies of the partonic structure of the nucleus using charm- and bottom-quarks production, as well as measurements highly relevant to astrophysics. For the research considered in this paper, the key advantage is the ability to easily change the ion species of the target, and therefore to vary the mass number of the target nucleus over a wide range, from a single proton to heavy elements. Additionally, these collisions are expected to produce large yields of charmed mesons, ensuring a good statistical precision for studies involving charm quarks. Furthermore, installing a polarised target would create unprecedented opportunities for studying the spin structure of a nucleon, which is a primary objective for the LHCspin initiative Santimaria:2021zay; Pappalardo:2024jpe.
@@
The LHCb detector LHCb:2008vvz is a single-arm spectrometer that provides track reconstruction, charge hadron and muon identification over a wide range of momentum, and a vertex detector allowing for efficient studies of charmed and bottom hadrons using their full reconstruction down to transverse momentum $P_{T}\approx 0$. It covers the forward rapidity region with geometrical acceptance in pseudorapidity of $52<\eta<$ 5. At LHCb, measurement of $D$ and $B$ mesons allows one to access relatively large values of $x_{2}$, the momentum fraction of the parton in the target: $0.06\lesssim x_{2}\lesssim 0.76$ for $B$ mesons and $0.02\lesssim x_{2}\lesssim 0.27$ in the case of $D$ mesons. [^1] The value of $x_{2}$ is estimated assuming $2\to 2$ kinematics for the $D$ or $B$ meson production process, which leads to a simple formula: $x_{1,2}=Me^{\pm y_{\rm cms}}/\sqrt{s}$, where $M$ is the mass of the meson. Additional estimates of the probed $x$ region based on NLO calculations are provided in Fig. 1. Such values of $x_{2}$ cover the part of EMC and antishadowing regions in the target nucleus, which are difficult to probe effectively in collider experiments.
@@
LHCb had successfully implemented the first fixed-target program using its System for Measuring Overlap with Gas (SMOG), which was developed for luminosity measurements. In this case, a noble gas was injected into the vertex locator (VELO) system, and the gas served as a target. LHCb collected samples of fixed-target collision data using the LHCb proton and lead beams during the LHC Run 2, performed their analysis and delivered important physics results LHCb:2024vwi; LHCb:2022qvj; LHCb:2022sxs; LHCb:2022cul; LHCb:2022bbb; LHCb:2018jry. The SMOG system provided an excellent demonstration of the feasibility of such studies, but it also had its limitations. The available luminosity was limited by the possible density of gas inside the VELO, and only a few ion species (He, Ne and Ar) were available for experimentation. These shortcomings were addressed with the new system, SMOG2, and the installation of a gas storage cell before the start of LHC Run 3 BoenteGarcia:2024kba. The storage cell allows one to inject gas into a well-defined volume within the LHC beam line. It facilitates a higher density of gas within the cell while still maintaining outside the vacuum level required for LHC operations. This upgrade increases the available luminosity in the fixed-target mode by up to two orders of magnitude, allowing for collecting up to 100 pb-1 data annually with a proton beam on the H2 target and enabling the use of new ion species, provided that the given gas is compatible with LHC operation. These improvements open new perspectives for measurements of charm- and bottom-quark production in $p$A collisions over a wide range of target atomic mass numbers A with unprecedented precision. In turn, it provides a unique opportunity to significantly improve the current understanding of gluon nPDFs.
II.2 Experimental conditions with SMOG2 at LHCb
@@
The target gas species and luminosities for SMOG2 that are used in this study were taken from Bursche:2018orf and are summarised in Table 1. These numbers were proposed as a possible data-taking scenario for Run 3. SMOG2 has already performed runs with H2, He, Ne and Ar target. Injections into SMOG2 of gases such as D2, N2, O2, Kr and Xe are being studied. In our study, that focuses on the atomic mass number dependence of nPDFs, D2 was not considered but it is also of great interest for the nPDFs extraction, in particular to test the isospin symmetry assumption of the nPDFs.
@@
Table 1: Summary of the possible collision systems and luminosities for the SMOG2 experiment for the LHC Run 3 Bursche:2018orf at $\sqrt{s_{{}_{NN}}}=113$ GeV. These luminosities are used for $D$ meson pseudo-data generation. The luminosities in parenthesis correspond to those used for $B$ meson pseudo-data generation as proposed in Ref. Hadjidakis:2018ifr.
Since the cross sections for $B$ mesons are small, higher luminosities indicated in parenthesis in Table 1 are also considered in this study in the case of H2 and Xe targets, as proposed in Ref. Hadjidakis:2018ifr. The rapidity acceptance of the LHCb detectors in the laboratory frame for open heavy-flavor measurements is taken as $4.62<y_{\rm lab}<$ 4.6, which translates to a center-of-mass (cms) coverage of $-2.79<y_{\rm cms}<-0.19$, corresponding to $\Delta y\simeq$ 4.79, for the collisions at $\sqrt{s_{NN}}=113$ GeV with a proton beam energy $E_{p}=6.8$ TeV. We consider studies with the $D^{0}$ meson reconstructed via its hadronic decay $D^{0}\to K^{-}\pi^{+}$, and $B^{+}$ measured through its decay into $D$ meson: $B^{+}\to D^{0}+\pi^{+}$ (and charge conjugate particles $\overline{D}^{0}$ and $B^{-}$, respectively). There is a possibility to improve the statistical precision for $B$-meson studies if other decay channels are considered, but combining such measurements with different systematic uncertainties is not always straightforward. Thus, we focus here on a single $B^{\pm}$ decay channel. The overall efficiency for $D$ mesons, that is the product of the geometrical acceptance of the detector and the efficiencies of the reconstruction, the selection, the PID and trigger requirements, is inspired by the SMOG analysis LHCb:2022qvj and from SMOG2 simulations for $D$ mesons for the rapidity dependence Santimaria:private. The integrated efficiency over rapidity is taken as 1% and it varies as a function of rapidity: from 0.04-0.3% at the lowest rapidities, reaching a maximum value of 2.1% at $y_{\rm lab}\approx 3.7$ and decreasing to 0.7% at $y_{\rm lab}\approx 4.6$. For the $B$ meson, there are not yet studies with the SMOG2 system and we assume a similar overall efficiencies as for the $D$ meson. We also assume the cross sections measurements will have systematic uncertainties of 4-5% for $D$ mesons and 8% for $B$ mesons in addition to the statistical ones. Those values are based on previous LHCb measurements in collider mode with large data samples LHCb:2017yua; LHCb:2019avm and the systematic uncertainties expectations on $J/\psi$ cross section measurements with SMOG2 of the order of 2-3% Bursche:2018orf.
III Methodology
III.1 Overview of the methodology
@@
    Computation of the cross section for charm- and bottom-quark production in $p$A reaction at $\sqrt{s_{NN}}$ = 113 GeV using the MadGraph5_aMC@NLO framework Flore:2025ync.
@@
III.2 Cross section for charm- and bottom-quark production
We calculate the cross section for charm- and bottom-quark production, $\sigma_{p{\rm A}\to c\overline{c}}$ and $\sigma_{p{\rm A}\to b\overline{b}}$, respectively, using the MadGraph5_aMC@NLO framework Alwall:2014hca (MG5aMC in what follows) extended to accommodate asymmetric reactions Flore:2025ync. MG5aMC allows for automated perturbative QCD calculations for standard model processes at Next-to-Leading Order (NLO) accuracy. The differential cross section $d\sigma_{p{\rm A}\to Q\overline{Q}X}$ (where $Q=c$ or $b$ quark) is given by the general formula:
@@ III.2 Cross section for charm- and bottom-quark production
where $A_{\rm A}$ is the atomic mass number of the target nucleus A, $f_{i/p}$ and $f_{j/{\rm A}}$ are the PDFs of the proton and the target nucleus A, $x_{i}$ and $x_{j}$ are the fractions of the longitudinal momentum carried by the partons compared to the proton and nucleus A, respectively; $\mu_{F}$ and $\mu_{R}$ are the factorisation and renormalisation scales, and $d\widehat{\sigma}_{ij\to Q\overline{Q}}$ is the parton-level cross section for production of heavy quark pair. The LHAID_p and LHAID_A indices indicate that the inputs are taken from the LHAPDF library Buckley:2014ana.
@@
$R_{p{\rm A}}=\frac{1}{A_{\rm A}}\frac{d\sigma_{p{\rm A}}}{d\sigma_{pp}},$ (2)
$R_{p{\rm A}}$ is very useful for nPDFs studies since theoretical and experimental uncertainties are expected to cancel to some extent within the ratio. Moreover, the MG5aMC version with asymmetric collisions Flore:2025ync provides both single cross sections for the $pp$ and $p$A collisions and nuclear modification factors themselves.
In our study, we use a selection of different nPDFs: nCTEQ15HIX (a recent update of the nCTEQ15 Kovarik:2015cma fit concentrating on the high-$x$ region) Segarra:2020gtj, nCTEQ15HQ (the newest nCTEQ PDF set that includes many heavy-flavor data from $p$Pb collisions at the LHC) Duwentaster:2022kpv, and EPPS21 Eskola:2021nhw (the latest set from EPPS group including data for $D$-meson production from $p$Pb collisions at the LHC). For the proton PDFs, we adopt the corresponding free-proton baseline PDFs Owens:2007kp; Hou:2019qau.
III.3 Pseudo-data generation

### v2

# The impact of future $D$- and $B$-meson measurements with the SMOG2 program at LHCb on the determination of nuclear parton distribution functions

@@ Abstract
We perform an analysis of the potential impact of future $D$- and $B$-meson measurement within the SMOG2 fixed-target program at the LHCb experiment on nuclear parton distribution functions. Following [1], we assume that SMOG2 will collect data for five nuclear targets: He, Ne, Ar, Kr, Xe and hydrogen which will provide a baseline for constructing nuclear ratios. The analysis is performed by using the PDF reweighting method. We demonstrate that such a measurement will allow to considerably reduce the current uncertainties of the nuclear gluon distribution and, to some extent, even the uncertainties of the light sea-quark distributions. Furthermore, it will provide the first possibility to systematically study the current assumptions on the $A$-dependence of the nuclear parton distributions.
keywords:
LHC , fixed-target experiment , SMOG2 , heavy flavour , nPDFs , reweighting
journal: Physics Letters B
[…]
[addr1]organization=Dipartimento di Fisica, Università di Cagliari, addressline=Cittadella Universitaria, city=Cagliari, postcode=I-09042, state=Monserrato (CA), country=Italy
[…]
[addr2]organization=INFN, Sezione di Cagliari, addressline=Cittadella Universitaria, city=Cagliari, postcode=I-09042, state=Monserrato (CA), country=Italy
[…]
[addr3]organization=Université Paris-Saclay, CNRS/IN2P3, IJCLab, ( city=Orsay, postcode=91405, country=France
[…]
[addr4]organization=Faculty of Physics, Warsaw University of Technology, addressline=plac Politechniki 1, city=Warszawa, postcode=00-661, country=Poland
[…]
[addr5]organization=Institute of Nuclear Physics, Polish Academy of Sciences, addressline=ul. Radzikowskiego 152, city=Cracow, postcode=31-342, country=Poland
1 Introduction
@@
One of the crucial aspects of high-energy nuclear physics is studying the structure of matter, in particular, in terms of its smallest components, quarks and gluons, collectively referred to as partons. Within the collinear factorisation approach [2, 3], the partonic structure of a hadron is described by parton distribution functions (PDFs), which represent the probability of finding partons with a given fraction $x$ of the proton longitudinal momentum. PDFs encapsulate the non-perturbative information about hadron content, and are determined through fits to experimental data [4, 5, 6, 7].
@@
Furthermore, the partonic structure of nucleons bound in a nucleus differs from that of a free proton. This phenomenon was first identified as the EMC effect [8]. Subsequent studies have shown that this modification varies with $x$ and we can now recognise four characteristic regions of nuclear modifications [9, 10, 11, 12, 13, 14, 15]. These are identified as: (i) the shadowing region at $x\lesssim 0.1$ (characterised by a depletion of nuclear cross-sections compared to the free-nucleon ones), (ii) the antishadowing region at $0.25\lesssim x\lesssim 0.1$ (characterised by a corresponding enhancement), (iii) the EMC region $0.75\lesssim x\lesssim 0.25$ (characterised by depletion), and (iv) the Fermi-motion region $x\gtrsim 0.75$ (characterised by enhancement). The strength of these effects also changes with the atomic mass number of the nucleus.
@@
Factorisation framework directly links these modifications observed on the level of cross sections with the behaviour of nuclear PDFs (nPDFs). Therefore, a precise determination of nPDFs is essential for using perturbative quantum chromodynamics (pQCD) to interpret experimental data from high-energy nuclear reactions. For example, this information is crucial for using charm and bottom mesons to study the hot and dense partonic matter created in ultrarelativistic heavy-ion collisions at the Relativistic Heavy Ion Collider at Brookhaven National Laboratory [16] or at the Large Hadron Collider (LHC) at CERN (for example [17]), as well as for computing cross sections in astrophysics, e.g. [18]. Despite significant progress over the past decades, our understanding of nPDFs, especially the gluon ones, remains incomplete. For instance, there is a limited knowledge about how nuclear effects vary with the number of nucleons in a nucleus.
@@
Charm and bottom quarks in high-energy reactions are predominantly produced via gluon fusion. Therefore, the yields of charm and bottom hadrons can be used to determine gluon nPDFs [19, 20]. The LHCb experiment facilitates an effective measurement of $D$ and $B$ mesons in proton-nucleus collisions across various ion species making it a perfect place for robust study on nuclear gluon PDFs.
@@
In this paper, we review the prospects for investigating nuclear parton distribution functions using $p$A pseudo-data at $\sqrt{s_{NN}}=113$ GeV from the SMOG2 fixed-target program at the LHCb experiment [21]. The pseudo-data used for the analysis are generated with MadGraph5_aMC@NLO [22]. The paper is organised as follows: in Section 2 we begin with a brief overview of the main features of fixed-target collisions using LHC beams. Next, we describe the experimental setup for these collisions at the LHCb experiment, including the ion species that can be used and the expected parameters of this experimental program. Section 3 introduces the formalism employed to evaluate the impact of the anticipated LHCb data on nPDFs. Finally, we present our results in Section 4 and summarise our findings in Section 5.
2 Fixed-target program at the LHCb experiment
2.1 Overview of the fixed-target program at the LHC
@@
An experiment utilising the LHC beams and a fixed target offers numerous, exciting and timely opportunities [23, 24]. The high energy and intensity of the proton and ion beams at the LHC enable compelling studies of the partonic structure of the nucleus using charm- and bottom-quarks production, as well as measurements highly relevant to astrophysics. For the research considered in this paper, the key advantage is the ability to easily change the ion species of the target, and therefore to vary the mass number of the target nucleus over a wide range, from a single proton to heavy elements. Additionally, these collisions are expected to produce large yields of charmed mesons, ensuring a good statistical precision for studies involving charm quarks. Furthermore, installing a polarised target would create unprecedented opportunities for studying the spin structure of a nucleon, which is a primary objective for the LHCspin initiative [25, 26].
@@
The LHCb detector [27] is a single-arm spectrometer that provides track reconstruction, charge hadron and muon identification over a wide range of momentum, and a vertex detector allowing for efficient studies of charmed and bottom hadrons using their full reconstruction down to transverse momentum $P_{T}\approx 0$. It covers the forward rapidity region with geometrical acceptance in pseudorapidity of $52<\eta<$ 5. At LHCb, measurement of $D$ and $B$ mesons allows one to access relatively large values of $x_{2}$, the momentum fraction of the parton in the target: $0.06\lesssim x_{2}\lesssim 0.76$ for $B$ mesons and $0.02\lesssim x_{2}\lesssim 0.27$ in the case of $D$ mesons. [^1] The value of $x_{2}$ is estimated assuming $2\to 2$ kinematics for the $D$ or $B$ meson production process, which leads to a simple formula: $x_{1,2}=Me^{\pm y_{\rm cms}}/\sqrt{s}$, where $M$ is the mass of the meson. Additional estimates of the probed $x$ region based on NLO calculations are provided in Fig. 1. Such values of $x_{2}$ cover the part of EMC and antishadowing regions in the target nucleus, which are difficult to probe effectively in collider experiments.
@@
LHCb had successfully implemented the first fixed-target program using its System for Measuring Overlap with Gas (SMOG), which was developed for luminosity measurements. In this case, a noble gas was injected into the vertex locator (VELO) system, and the gas served as a target. LHCb collected samples of fixed-target collision data using the LHCb proton and lead beams during the LHC Run 2, performed their analysis and delivered important physics results [28, 29, 30, 31, 32, 33, 34, 35]. The SMOG system provided an excellent demonstration of the feasibility of such studies, but it also had its limitations. The available luminosity was limited by the possible density of gas inside the VELO, and only a few ion species (He, Ne and Ar) were available for experimentation. These shortcomings were addressed with the new system, SMOG2, and the installation of a gas storage cell before the start of LHC Run 3 [21]. The storage cell allows one to inject gas into a well-defined volume within the LHC beam line. It facilitates a higher density of gas within the cell while still maintaining outside the vacuum level required for LHC operations. This upgrade increases the available luminosity in the fixed-target mode by up to two orders of magnitude, allowing for collecting up to 100 pb-1 data annually with a proton beam on the H2 target and enabling the use of new ion species, provided that the given gas is compatible with LHC operation. These improvements open new perspectives for measurements of charm- and bottom-quark production in $p$A collisions over a wide range of target atomic mass numbers A with unprecedented precision. In turn, it provides a unique opportunity to significantly improve the current understanding of gluon nPDFs.
2.2 Experimental conditions with SMOG2 at LHCb
@@
The target gas species and luminosities for SMOG2 that are used in this study were taken from [1] and are summarised in Table 1. These numbers were proposed as a possible data-taking scenario for Run 3. SMOG2 has already performed runs with H2, He, Ne and Ar target. Injections into SMOG2 of gases such as D2, N2, O2, Kr and Xe are being studied. [^2] Recent measurements with a D2 target were reported at the Quark Matter conference [36]. In our study, that focuses on the atomic mass number dependence of nPDFs, D2 was not considered but it is also of great interest for the nPDFs extraction, in particular to test the isospin symmetry assumption of the nPDFs.
@@
Table 1: Summary of the possible collision systems and luminosities for the SMOG2 experiment for the LHC Run 3 [1] at $\sqrt{s_{{}_{NN}}}=113$ GeV. These luminosities are used for $D$ meson pseudo-data generation. The luminosities in parenthesis correspond to those used for $B$ meson pseudo-data generation as proposed in Ref. [23].
Since the cross sections for $B$ mesons are small, higher luminosities indicated in parenthesis in Table 1 are also considered in this study in the case of H2 and Xe targets, as proposed in Ref. [23]. The rapidity acceptance of the LHCb detectors in the laboratory frame for open heavy-flavor measurements is taken as $4.62<y_{\rm lab}<$ 4.6 [29], which translates to a center-of-mass (cms) coverage of $-2.79<y_{\rm cms}<-0.19$, corresponding to $\Delta y\simeq$ 4.79, for the collisions at $\sqrt{s_{NN}}=113$ GeV with a proton beam energy $E_{p}=6.8$ TeV. We consider studies with the $D^{0}$ meson reconstructed via its hadronic decay $D^{0}\to K^{-}\pi^{+}$, and $B^{+}$ measured through its decay into $D$ meson: $B^{+}\to D^{0}+\pi^{+}$ (and charge conjugate particles $\overline{D}^{0}$ and $B^{-}$, respectively). There is a possibility to improve the statistical precision for $B$-meson studies if other decay channels are considered, but combining such measurements with different systematic uncertainties is not always straightforward. Thus, we focus here on a single $B^{\pm}$ decay channel. The overall efficiency for $D$ mesons, that is the product of the geometrical acceptance of the detector and the efficiencies of the reconstruction, the selection, the PID and trigger requirements, is inspired by the SMOG analysis [33] and from SMOG2 simulations for $D$ mesons for the rapidity dependence [37]. The integrated efficiency over rapidity is taken as 1% and it varies as a function of rapidity: from 0.04-0.3% at the lowest rapidities, reaching a maximum value of 2.1% at $y_{\rm lab}\approx 3.7$ and decreasing to 0.7% at $y_{\rm lab}\approx 4.6$. For the $B$ meson, there are not yet studies with the SMOG2 system and we assume a similar overall efficiencies as for the $D$ meson. We also assume the cross sections measurements will have systematic uncertainties of 4-5% for $D$ mesons and 8% for $B$ mesons in addition to the statistical ones. Those values are based on previous LHCb measurements in collider mode with large data samples [38, 39] and the systematic uncertainties expectations on $J/\psi$ cross section measurements with SMOG2 of the order of 2-3% [1]. In addition, we assume the systematic uncertainties are uncorrelated between the $pp$ and $p{\rm A}$ collision systems.
3 Methodology
3.1 Overview of the methodology
@@
    Computation of the cross section for charm- and bottom-quark production in $p$A reaction at $\sqrt{s_{NN}}$ = 113 GeV using the MadGraph5_aMC@NLO framework [22].
@@
3.2 Cross section for charm- and bottom-quark production
We calculate the cross section for charm- and bottom-quark production, $\sigma_{p{\rm A}\to c\overline{c}}$ and $\sigma_{p{\rm A}\to b\overline{b}}$, respectively, using the MadGraph5_aMC@NLO framework [40] (MG5aMC in what follows) extended to accommodate asymmetric reactions [22]. MG5aMC allows for automated perturbative QCD calculations for standard model processes at Next-to-Leading Order (NLO) accuracy. The differential cross section $d\sigma_{p{\rm A}\to Q\overline{Q}X}$ (where $Q=c$ or $b$ quark) is given by the general formula:
@@ 3.2 Cross section for charm- and bottom-quark production
where $A_{\rm A}$ is the atomic mass number of the target nucleus A, $f_{i/p}$ and $f_{j/{\rm A}}$ are the PDFs of the proton and the target nucleus A, $x_{i}$ and $x_{j}$ are the fractions of the longitudinal momentum carried by the partons compared to the proton and nucleus A, respectively; $\mu_{F}$ and $\mu_{R}$ are the factorisation and renormalisation scales, and $d\widehat{\sigma}_{ij\to Q\overline{Q}}$ is the parton-level cross section for production of heavy quark pair. The LHAID_p and LHAID_A indices indicate that the inputs are taken from the LHAPDF library [41].
@@
$R_{pA}=\frac{1}{A_{A}}\frac{d\sigma_{pA}}{d\sigma_{pp}},$ (2)
$R_{pA}$ is very useful for nPDFs studies since theoretical and experimental uncertainties are expected to cancel to some extent within the ratio. Moreover, the MG5aMC version with asymmetric collisions [22] provides both single cross sections for the $pp$ and $p$A collisions and nuclear modification factors themselves.
In our study, we use a selection of different nPDFs: nCTEQ15HIX (a recent update of the nCTEQ15 [42] fit concentrating on the high-$x$ region) [43], nCTEQ15HQ (the newest nCTEQ PDF set that includes many heavy-flavor data from $p$Pb collisions at the LHC) [44], and EPPS21 [45] (the latest set from EPPS group including data for $D$-meson production from $p$Pb collisions at the LHC). For the proton PDFs, we adopt the corresponding free-proton baseline PDFs [46, 47].
3.3 Pseudo-data generation